Course: Theory of Interest

FM) A perpetuity is purchased for $7,000. It's first annual payment of $200 will occur five years from now. Each subsequent payment is increased by an amount C from the previous payment (the payments as 200, 200 + C, 200 + 2C, ...). If the effective annual interest rate is 4% find the value of C.

Answer: $4.3615664730
Answer #1

perpetuity = annual cash flow / discount rate- g

7000 = 200 / 4% - g

   7000 g - 280 = 200

g = 80 / 7000

g = 1.142857

note: where g stand for growing rate